OpenStudy (anonymous):

SOMEONE PLEASE HELP! A rectangular yard measuring 40 ft by 71ft is bordered (and surrounded) by a fence. Inside, a walk that is 3 ft wide goes all the way along the fence. Find the area of this walk. Be sure to include the correct unit in your answer.

OpenStudy (anonymous):

is 40ft the length and 71ft the width of the yard?

OpenStudy (anonymous):

i'm guessing. that is how it is written. so yes.

OpenStudy (anonymous):

it is worded really weird

OpenStudy (anonymous):

ok then, if 40ft is the length, u can multiply that times the width of the walk, so 3 ft, 40*3= 120...i think that would be the area 120 sq ft

OpenStudy (anonymous):

it said that was incorrect

OpenStudy (anonymous):

it will show me how to do it. and iwill tell u. then will u heelp me with the next 2 problems like this one?

OpenStudy (anonymous):

sure ill try to help u

OpenStudy (anonymous):

We start by drawing a rectangle representing the fence surrounding the yard. Inside the fence, we draw a wide walk going along the fence. inner rectangle's length is the yard length twice the walk width. so 71-3-3=65 ft and the width is 40-3-3=34 ft

OpenStudy (anonymous):

since the walk is the region between the 2 rectangles..it's area is the area of the larger rectangle minus the area of the smaller rectangle. so a=40x71-34x65 =2840-2210 =630 ft ^2

OpenStudy (anonymous):

follow? bc it's kinda confusing to me.

OpenStudy (anonymous):

yes, i think i follow

OpenStudy (anonymous):

ok, i got it :)

OpenStudy (anonymous):

it sounds more like a rectangular box?

OpenStudy (anonymous):

i know. this thing is so stupid..but i think i have the hang of it now.

OpenStudy (radar):

It kinda makes sense, otherwise you might include the corners twice.

OpenStudy (radar):

Did you work the next one?